OpenStudy (anonymous):

A 30 kg block moving initially slideing on a level surface at 15 m/s is subject to a constant friction of 5N untill it comes to a stop. How far will it slide and how much energy is dissipated due to the friction

OpenStudy (matt101):

This is the exact same question you asked before, just with different numbers. See my answer there and just plug in the numbers you have here. http://openstudy.com/study#/updates/5508d6afe4b015f1766da28a

OpenStudy (anonymous):

I got 3309J

OpenStudy (anonymous):

Would you say that's right

OpenStudy (matt101):

Show me your calculations

OpenStudy (anonymous):

F=ma so -5=30a and a=.17, So 0 = 225-2(.17) d=661.8, w=(5)(661.8) so w= 3309

OpenStudy (matt101):

Looks good! I just didn't round the acceleration and kept it as 1/6 so I came up with 3375 instead.